# Smart walkers in discrete space

### 12 Mar 2026 v2

@@ 1 Introduction
The study of random walkers has historically provided invaluable insights into a wide range of complex phenomena. Examples can be found in statistical physics [19], network theory [13, 8, 12], robotics [24] or finance [5, 4]. Random walkers often navigate in discrete spaces by using simple stochastic rules. Therefore, they offer a reference baseline to elaborate on a large plethora of complex dynamical schemes. These latter may account for adaptability, memory, or even strategic interactions to enable for a comprehensive description of the different phenomena under scrutiny.
@@
The manuscript is structured as follows: In Sec. 2 we describe the model of two random walkers on a Markov chain as a game. In Sec. 3, we present the simplest case of two random walkers. We provide an analytical derivation of the probability distribution of the first encounter as a new result, and then we re-derive known results [14] through a novel and simpler proof detailed in Appendix A. This result constitutes the benchmark reference for the subsequent analysis. In Sec. 4, we illustrate how to incorporate a smart strategy into a walker, outlining the reinforcement learning training procedure. We also introduce the information-theoretic and thermodynamic metrics utilized in our analysis. Sec. 4.2 presents our numerical results, achieved by training a random walker using linear, sinusoidal, and non-Markovian reward policies. Moreover, in Sec. 5, and motivated by the success of the above analysis, we speculate that configuration entropy can be employed as an effective measure of attitude or intelligent behaviour. To elaborate in this direction we also consider a specific application to table top games. Finally, we conclude in Sec. 6 by discussing the limitations and strengths of our research. The case of two simultaneously trained agents ( interacting via a predator- prey scheme) is shortly addressed in the annexed Appendix C to show the versatility of the proposed approach, beyond the main setting here addressed.
@@ 4.2 The imprint of learning on the recorded statistics
The reward profile (4.9) was chosen as it resembles the natural encounter profile of completely random, untrained, agents. The other two linear profiles were instead selected to sketch a toy zero-sum scenario, with or without time pressure. One could interpret, for example, the two walkers as a buyer-seller couple, moving in price space during trading, as already mentioned in the introductory section. The agent (Alice) was trained for 10,00010{,}000 episodes under each reward function, using an inverse temperature parameter $\beta$ that started at an initial value of $1$ and was capped at an upper threshold of $10$. This training regime yielded dynamical systems of the chaser-target typology with substantially morphed first-encounter probability distributions (see Figure 6) and mean first-encounter times (see Figure 7). In the aforementioned Figures, the histogram refer to actual simulation of the learned stochastic dynamics. The solid (red) lines stand for the analytical solutions, computed following the procedure introduced in Section 3, and by making use of the policy table to which Alice converged upon learning, subject to the different assumed rewards.
@@ 6 Conclusions
Summing up, working in the framework of a chaser-target model, with (just one) trainable agent, and leveraging on established reinforcement learning strategies, we proposed and tested the notion of configurational entropy as a possible measure of the rational - non random - operating skills. Extending this analysis to a setting where the walkers are both allowed to learn while exploring a more intricate spatial environment (a lattice in higher dimensions or a heterogeneous graph) constitutes a future direction of possible investigation.

### 30 Mar 2026 v3

@@ 1 Introduction
The study of random walkers has historically provided invaluable insights into a wide range of complex phenomena. Examples can be found in statistical physics [21, 2, 1, 23], network theory [15, 10, 14, 24], robotics [28] or finance [7, 6]. Random walkers often navigate in discrete spaces by using simple stochastic rules. Therefore, they offer a reference baseline to elaborate on a large plethora of complex dynamical schemes. These latter may account for adaptability, memory, or even strategic interactions to enable for a comprehensive description of the different phenomena under scrutiny.
@@
The manuscript is structured as follows: In Sec. 2 we describe the model of two random walkers on a Markov chain as a game. In Sec. 3, we present the simplest case of two random walkers. We provide an analytical derivation of the probability distribution of the first encounter as a new result, and then we re-derive known results [16] through a novel and simpler proof detailed in Appendix A. This result constitutes the benchmark reference for the subsequent analysis. In Sec. 4, we illustrate how to incorporate a smart strategy into a walker, outlining the reinforcement learning training procedure. Specifically, one of the two agents is adaptively trained to eventually develop an intelligent walk, under a reward that is being set a priori. At the end of the process, the smart walker has learned a strategy which is then applied during the testing phase: at each iteration the intelligent walker takes a move that is not dependent on past history. It only depends on its own position and on the position of the opponent, which keeps on performing an unbiased random walk. The pattern of possible moves is hence encoded in a generalized states space (the smart walker acts differently depending on its own position and on the position occupied by the opponent, at the time the move has to be taken). The dynamics is therefore exactly Markovian, namely it follows a stochastic process where the future state depends solely on the current state - in this case, of both walkers - not on the past history. This enables us to generalize the above formal analysis to the relevant setting where one (potentially two) walker(s) evolve following an intelligent strategic plan. The information-theoretic and thermodynamic metrics utilized in our analysis is also introduced. Sec. 4.2 presents our numerical results, achieved by training a random walker using linear, sinusoidal, and time-dependent reward policies. Moreover, in Sec. 5, and motivated by the success of the above analysis, we speculate that configuration entropy can be employed as an effective measure of attitude or intelligent behaviour. To elaborate further in this direction we consider a real world application to table top games. Finally, we conclude in Sec. 6 by discussing the limitations and strengths of our research. The case of two simultaneously trained agents ( interacting via a predator- prey scheme) is shortly addressed in the annexed Appendix C to show the versatility of the proposed approach, beyond the main setting here addressed.
@@ 4.2 The imprint of learning on the recorded statistics
The reward profile (4.9) was chosen as it resembles the natural encounter profile of completely random, untrained, agents. The other two linear profiles were instead selected to sketch a toy zero-sum scenario, with or without time pressure. One could interpret, for example, the two walkers as a buyer-seller couple, moving in price space during trading, as already mentioned in the introductory section. The agent (Alice) was trained for 10,00010{,}000 episodes under each reward function, using an inverse temperature parameter $\beta$ that started at an initial value of $1$ and was capped at an upper threshold of $10$. This training regime yielded dynamical systems of the chaser-target typology with substantially morphed first-encounter probability distributions (see Figure 6) and mean first-encounter times (see Figure 7). In the aforementioned Figures, the histogram refer to actual simulation of the learned stochastic dynamics. The solid (red) lines stand for the analytical solutions, computed following the procedure introduced in Section 3, and by making use of the policy table to which Alice converged upon learning, subject to the different assumed rewards. A comment is mandatory at this point. As also anticipated in the introduction, the problem of a smart walker chasing (post training) a target opponent (which still performs pure random walk) is intrinsically Markovian, when examined in the expanded space of the $N^{2}$ states for the co-evolving pair. No approximations are involved in the formal analysis reported above, which is therefore exact.
@@ 6 Conclusions
Summing up, working in the framework of a chaser-target model, with (just one) trainable agent, and leveraging on established reinforcement learning strategies, we proposed and tested the notion of configurational entropy as a possible measure of the rational - non random - operating skills. This is the main message that we intend to deliver. Indeed, the configurational entropy has been validated as viable quantity to measure the level of awareness acquired by Alice upon training. This latter quantity correlates nicely with the policy entropy, which cannot be straightforwardly accessed in real world experiments, as we argued above. Nicely enough, we show that the configurational entropy can be successfully applied to a real world challenge (a chess game) for which the policy entropy is not directly available. Configurational entropy is thus a solid measure of agents’ awareness. The investigated toy model (the two walkers sharing a discrete line) served as a benchmark framework to substantiate our claim, in a controlled setting that makes analytical progress possible (before and after training, in light of the inherent Markovianity of both models). Extending this analysis to a setting where the walkers are both allowed to learn while exploring a more intricate spatial environment (a lattice in higher dimensions or a heterogeneous graph) constitutes a future direction of possible investigation. For a first step in this direction, see Appendix C where two walkers interacting via a predator-prey scheme undergo simultaneous training.